OpenStudy (anonymous):

Which pair of equations are perpendicular lines? B.x – 3y = 11 and y = –3x + 23 C.4x + 2y = 6 and y = –2x + 18 D.x – 3y = 5 and y = 3x – 11

OpenStudy (texaschic101):

perpendicular lines will have negative reciprocal slopes. All that means is " flipping" the slope and changing the sign. For example : 1/3 and -3/1 or just -3 are perpendicular. (see how the slope is flipped and the sign is changed). So basically, in this problem you have to find the slopes and compare them. Do you know how to find the slope ?

OpenStudy (texaschic101):

To find the slope, put your equation in y = mx + b form, and the number in the m position is the slope. B. x - 3y = 11 --- subtract x from both sides -3y = -x + 11 -- now divide by -3 y = 1/3x - 11/3 the slope here is 1/3. Now the other equation is already in y = mx + b form. y = -3x + 23....slope on this one is -3. So our slopes are 1/3 and -3.....these are perpendicular lines.